Instructions For each of the following 4 problems provide 1. inputs 2. outputs 3. error conditions 4. an algorithm using pseudocode 5. the minimum and the maximum number of operations 6. a set of test cases 1. Multiplication Compute and display the result of how much a certain number of pounds of apple costs. 2. Gas Compute and display the price a person will pay for gas at the gas station. If the person pays with a credit card, there is an extra charge of 10% of the total price 3. University tuition Compute the tuition a student pays at an university, given the following rules: Every student pays a campus fee of $6.87 .Each one credit costs $20 An university employee pays only campus fee State employee pays campus fee + 10% of total number of credits cost Assume a student cannot be an university employee and a state employee at the same time A student paying in installments pays an additional 3% of total amount owed 4. Zoo census There are a number of zoos in the United States. Each of these zoos is a home to giraffes, lions and snakes. Determine and display the total number of each (giraffes, lions, and snakes) in the zoos in the United States. Assume positive values.
Answer #1

Solution:

1)

input: number of pounds of apple to be bought, cost of the apple per pound

Output: Expense to be made to buy apples.

Error condition: number of pounds <0, price per pound < 0

Algorithm using pseudo-code:

Step 1 : Declare variables Ans = 'Y', no_of_pounds, price_per_pound, cost

Step2 : While ( Ans = 'Y')

Step2.1 : Read no_of_pounds,price_per_pound

Step2.2 : cost := no_of_pounds * price_per_pound

Step2.3 : Print cost

Step2.4 : Provide the option for continue (Y) or abort (N)

Step2.5 : Read ans

  

Step3: End.

The minimum number of operation : 1 (As, one will go to the loop at least once,beacuse the value of 'ans' is 'Y' by default.

Maximum operations: As many times as one wants. As many times the user will want to continue the loop will iterate.

Set of test cases:

1.

no_of_pounds = 2

price_per_pound = 4

2.

no_of_pounds = 6

price_per_pound = 4

2)

Input: Litre of gas to be bought, Price per liter, Payment type

Output: Payment to be made

Error Condition: Litre of gas to be bought<0, Price per liter <0

Pseudo-code:

Step1: Declare variables Litre of gas to be bought, Price per liter, ans, Cost, choice = 1

Step2: while (choice == 1)

Step2.1: Read Litre of gas to be bought,  Price per liter

Step 2.2: Print " Do you have the credit card, Press Y for yes and N for no

Step 2.3: Read ans

Step 2.4: if (ans = 'Y')

Step 2.4.1: cost =  Litre of gas to be bought * Price per liter * 1.1

Step 2.4.2: Print cost

Step2.5: if (ans = 'N')

Step 2.5.1: cost =  Litre of gas to be bought * Price per liter

Step 2.5.2: Print cost

Step 2.6: Print "Do you want to continue? if yes enter 1

Step 2.7: Read choice

Step 3: Exit

Minimum No of Operations: 1 ( As the value of option = 1)

Maximum No of Operations: As many as you want

Set of test cases:

1. Litre of gas to be bought = 2

Price per liter, Payment type = 20

2. Litre of gas to be bought = 3

Price per liter, Payment type = 10

3)

3. Pseudo code :

  1. Set Campus Fee CF to $6.87
  2. Set Credit Costs CC to $20
  3. Read student employment status ES
  4. If ES is university employee

Then SET ES to “UE”

        Else if ES is State Employee

Then SET ES to “SE”

         Else

       Set ES to “NE”

  1. If ES equals to UE then set FEE=CF (i.e. $6.87)

Else if ES equals to SE then set FEE=(CF+10/100*CC)

Else set FEE=CF+CC

  1. Read Payment Type PT
  2. If PT is by installments, then set PT to INS else FULL
  3. If PT equal to INS then

FEE=FEE+3/100*FEE

Else

FEE

       I .   Write FEE

4. Pseudo code :

  1. Read the number of Zoo in united states ‘NZ’
  2. Set Total no of giraffes, lions, snakes i.e. TNG, TNL, TNS to 0
  3. For i in 1 to ‘NZ’

Read no of giraffes in zoo(i) i.e. NG

Read no of lions in zoo(i) i.e. NL

Read no of giraffes in snakes (i) i.e. NS

Set TNG=TNG+NG

Set TNL=TNL+NL

Set TNS=TNS+NS

End loop

4. Write TNG , TNL,TNS

I hope this helps if you find any problem. Please comment below. Don't forget to give a thumbs up if you liked it. :)